Find the radius of a curve banked at an angle of 35O negotiated by a car at a velocity of 20.0 m/s.


1
Expert's answer
2021-12-02T10:07:29-0500

The radius is


"r=\frac{v^2}{g\tan\theta}=\frac{20^2}{9.8\tan35°}=58.3\text{ m}."